OpenStudy (anonymous):

Solve the equations: 19 = h/3 - 8 0.6 + 0.8 = 1.4 PLEASE JUST ANSWER WITH THE ANSWER AND AN EXPLANATION

OpenStudy (anonymous):

The second equation has nothing to solve.

OpenStudy (anonymous):

Apparently it does? The question says: Solve the equation. Explain each step and identify the property used to reach each step

OpenStudy (anonymous):

Do you know anything about solving equations.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I really do not want to sound dumb but no I know nothing.

OpenStudy (anonymous):

\[ 19 = \frac h3 - 8 \]In this case we want to solve for \(h\). To do that we want to isolate it. That will leave us with: \[ h=\ldots \]Where \(\ldots\) is the value for \(h\). We can isolate \(h\) by doing operations to both sides.

OpenStudy (anonymous):

For example, the first thing we can try to get rid of is the \(8\).

OpenStudy (anonymous):

The opposite of subtracting \(8\) is to add \(8\).

OpenStudy (anonymous):

So we add \(8\) to both sides. \[ 19+8=\frac h3 -8 + 8 \]This is called the additive property of equality. It says we can add things to both sides and they will still be equal.

OpenStudy (anonymous):

Then we simplify:\[ 27=\frac h3 \]We know we have made progress because the \(h\) is getting closer to being isolated.

OpenStudy (anonymous):

We need to get rid of division by \(3\), and the opposite operation for this is to multiply by \(3\). \[ 27\times 3=\frac h3\times 3 \]This is called the multiplicative property of equality

OpenStudy (anonymous):

Again, we have to simplify: \[ 81 = h \]

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

Yeah kinda. i'm going to read over it a few times to make sure I can know it better. Thanks for the help. But what about the second one? My book says that it needs to be solved?

OpenStudy (anonymous):

\[ 0.6 + 0.8 = 1.4 \]For this one, you just do addition, to simplify:\[ 1.4=1.4 \]And so the answer is \(\text{true}\). There is no variable to solve.

OpenStudy (anonymous):

Okay thanks. Have a nice day =D